OpenStudy (anonymous):

MEDAL given work two problems out pleaseee 1. x^2- 8x -7=0 2. x^2 -3x=40

OpenStudy (anonymous):

1 .x²-8x-1 = 0 Use quadratic formula: a = 1, b = -8, c = -1 x = [-b±√(b²-4ac)]/(2a) x = {-(-8)±√[(-8)²-4(1)(-1)]}/[2(1)] x = {8±√[(64)+(4)]}/2 x = [8±√(68)]/2 x = [8±2√(17)]/2 x = 4±√(17) Solutions: x = 4±√(17)

OpenStudy (anonymous):

im sorry that's wrong @NateR

OpenStudy (anonymous):

K